# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v1

@@ Appendix C Channel transformations
We first describe a $ph$ to $pp$ channel transformation via the $\overline{ph}$ channel:
$p$ $h$$\longrightarrow$ $\overline{ph}$$\longrightarrow$ $pp$ (51)
@@
where we have to be a little careful with the arrows. The picture corresponds to the following, e.g. for the full vertex:
$F_{pp}(\nu,\nu^{\prime},\omega)$ $=F_{ph}(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}),$ (52)
$F_{\overline{ph}}(\nu,\nu^{\prime},\omega)$ $=F_{ph}(\nu,\nu+\omega,\nu^{\prime}-\nu),$ (53)
$F_{pp}(\nu,\nu^{\prime},\omega)$ $=F_{\overline{ph}}(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu).$ (54)
In the following, we will denote the “old” variables in every transformation step with a tilde. For the $ph\rightarrow\overline{ph}$ transformation
$F_{\overline{ph}}(\nu,\nu^{\prime},\omega)$ $=F_{ph}(\nu,\nu+\omega,\nu^{\prime}-\nu)=F_{ph}(\tilde{\nu},\tilde{\nu^{\prime}},\tilde{\omega})$ (55)
@@
with ($\textrm{with }(\tilde{\nu},\tilde{\nu^{\prime}},\tilde{\omega})=(\nu,\nu+\omega,\nu^{\prime}-\nu).$
We also need to shift the indices so that for example for $n=0$ and $m=0$ ($\nu+\omega=\frac{\pi}{\beta}=\tilde{\nu^{\prime}}$) we are at $\tilde{n^{\prime}}=0$ again.
$0\leq a,$ $b,c\leq N-1$
$n$ $=a-\frac{N}{2}$
$n^{\prime}$ $=b-\frac{N}{2}$
$m$ $=c-\frac{N}{2}$ (57)
@@
$\bar{a}$ $=a,\quad\textrm{no shift}$
$\bar{b}$ $=a+b-\frac{N}{2}=n+m+\frac{N}{2},\quad\textrm{shift by }\frac{N}{2}$
$\bar{c}$ $=-a+b+\frac{N}{2}=-n+n^{\prime}+\frac{N}{2},\quad\textrm{shift by}\frac{N}{2}.$ (58)
Hence, we get the shift vector $\bm{b}=(0,\frac{N}{2},\frac{N}{2})^{\mathrm{T}}$. For example, at $n=n^{\prime}$ we need $\omega=0$ and, thus, $\bar{c}=\frac{N}{2}$, which is ensured by the shift. This first transformation can be represented by an MPO with $D_{max}=9$.
For the $\overline{ph}\rightarrow pp$ transformation
$F_{pp}(\nu,\nu^{\prime},\omega)$ $=F_{\overline{ph}}(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu)=F_{\overline{ph}}(\tilde{\nu},\tilde{\nu^{\prime}},\tilde{\omega}),$ (59)
@@
with ($\textrm{with }(\tilde{\nu},\tilde{\nu^{\prime}},\tilde{\omega})=(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu).$
Using the same procedure as above, we get the shift vector $\bm{b}=(0,\frac{N}{2}-1,\frac{N}{2})^{\mathrm{T}}$. This affine transformation can be represented by an MPO with $D_{max}=15$.
@@
where we have to be a little careful with the arrows. The picture corresponds to the following, e.g. for the full vertex
$F_{ph}(\nu,\nu^{\prime},\omega)$ $=F_{pp}(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}),$ (62)
$F_{\overline{pp}}(\nu,\nu^{\prime},\omega)$ $=F_{pp}(\nu,\nu+\omega,\nu^{\prime}-\nu),$ (63)
$F_{ph}(\nu,\nu^{\prime},\omega)$ $=F_{\overline{pp}}(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu).$ (64)
@@ Appendix D Parquet equation in QTT format
Figure 14: (a) Absolute normalized error of the parquet equation with QTTs for $\Gamma_{d}$ compared to the exact values at $\omega=0$ in the fermionic Matsubara frequency plane for ${D_{\mathrm{max}}}=200$, $R=7$, $\beta=U=1$, $\epsilon=10^{-8}$. Runtime of the parquet equation for $\Gamma_{d}$ for various maximum bond dimensions and grid size parameters $R$ with $\beta=U=1$. The dashed lines indicate the cubic runtime increase with ${D_{\mathrm{max}}}$ (b) and linear increase by increasing R (c), which corresponds to an exponential increase in the number of grid points.
@@ Appendix E SDE in QTT format
Figure 15: Plot of the maximum absolute normalized error of the SDE with QTTs for the self-energy $\Sigma$ compared to the exact values for various grid sizes depending on the maximum bond dimension with $U=\beta=1$. The dashed lines indicate the results from the dense grid calculation without QTTs.
@@ Appendix F Technical limitations
In the calculations, a specified maximum bond dimension is set not only in the initial TCI, but also in every QTT operation, in order to avoid a blowing up of the bond dimensions since this is the computational bottleneck. As it was shown in Sec. V.1, for the initial TCI already bond dimensions slightly above 100 are sufficient to reach maximum normalized errors of $10^{-6}$ of the QTTs with respect to the exact data. However, a problem emerges in the QTT operations, which are at the moment SVD based and make use of the truncate function in ITensors.jl to compress the resulting QTTs back to a certain maximum bond dimension. Furthermore, the fit algorithm used for MPO-MPO contractions relies on the SVD truncation internally [39]. The SVD truncation minimizes the difference between an original MPS and an approximated one in terms of the Frobenius norm. Because the Frobenius norm of the vertex functions grows exponentially with $R$ due to a constant term, the SVD truncation is expected to fail at large $R$; the Frobenius norm reaches $c(2^{R})^{3}\approx c\times 3\times 10^{13}$ at $R=15$ ($c$ is the constant term).
Here, we have observed the truncation error to become more significant the larger the value of the grid parameter $R$ is and even leads to wrong results around $R=15$.
Figure 16: Absolute normalized error of the reconstructed $F_{d}$ compared to the exact data for the innermost $16\times 16$ fermionic Matsubara frequency grid at $\omega=0,\beta=U=1$. (a) and (b) show the error after using TCI with a set tolerance of $10^{-10}$ and the maximum bond dimensions set to 160 and 140 respectively. (c)-(f) show the error after applying the SVD based truncation to the QTT with maximum bond dimension 160 (a) with a set maximum bond dimension of 140. The truncation error increases with larger values of the grid parameter $R$.
@@
In Fig. 16, we show this behavior in case of the full vertex in the density channel $F_{d}$. In Figs. 16(a-b) the absolute normalized error is shown after applying TCI to evaluate a QTT for $F_{d}$ for different maximum bond dimensions. It can be seen that the error in this center ($16\times 16$) fermionic Matsubara frequency box is of $\mathcal{O}(10^{-15})$. Figures 16(c-f) show the error after applying the SVD based truncation to the QTT with maximum bond dimension 160 down to a maximum bond dimension of 140 for various grid sizes determined by $R$. Although using TCI with a maximum bond dimension of 140 [Fig. 16(b)] the QTT was able to reconstruct the exact data with an absolute normalized error of $\mathcal{O}(10^{-15})$, applying the SVD based truncation significantly worsens the results. Moreover, it can be seen that the error increases significantly with increasing $R$. This is why in the case of the iterative parquet solutions for the Hubbard atom, only results up to $R=11$ are shown, since the resulting maximum normalized error does not improve anymore for larger grids due to the truncation errors. This can also already seen in Fig. 11 for $R=11$, where the maximum normalized error at a maximum bond dimension of 200 is only slightly lower than in the case of $R=10$. However, this should only be a problem of the current implementation and first numerical tests indicate that it is possible to overcome this limitation in the future e.g. by using CI based truncation [32]. This is because the CI-based truncation relies on the maximum norm and thus does not suffer from the divergence of the Frobenius norm.

### v2

@@ Appendix C Channel transformations
We first describe a ph $\rightarrow$ pp channel transformation via the $\overline{\textit{ph}}$ channel:
[…]
@@
where $\nu^{(^{\prime})}=(2n^{(^{\prime})}+1)\pi/\beta$ and $\omega=2m\pi/\beta$. The picture corresponds to the following, e.g. for the full vertex:
$F_{\textit{pp}}(\nu,\nu^{\prime},\omega)$ $=F_{\textit{ph}}(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}),$ (52a)
$F_{\overline{\textit{ph}}}(\nu,\nu^{\prime},\omega)$ $=F_{\textit{ph}}(\nu,\nu+\omega,\nu^{\prime}-\nu),$ (52b)
$F_{\textit{pp}}(\nu,\nu^{\prime},\omega)$ $=F_{\overline{\textit{ph}}}(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu).$ (52c)
In the following, we will denote the “old” variables in every transformation step with a tilde. For the ph to $\overline{\textit{ph}}$  transformation
$F_{\overline{\textit{ph}}}(\nu,\nu^{\prime},\omega)$ $=F_{\textit{ph}}(\nu,\nu+\omega,\nu^{\prime}-\nu)=F_{\textit{ph}}(\tilde{\nu},\tilde{\nu}^{\prime},\tilde{\omega})$ (53)
@@
with $(\tilde{\nu},\tilde{\nu}^{\prime},\tilde{\omega})=(\nu,\nu+\omega,\nu^{\prime}-\nu)$ and $\tilde{\nu}^{(^{\prime})}=(2\tilde{n}^{(^{\prime})}+1)\pi/\beta$ and $\tilde{\omega}=2\tilde{m}\pi/\beta$. We also need to shift the indices such that, for example for $n=0$ and $m=0$ ($\nu+\omega=\frac{\pi}{\beta}=\tilde{\nu}^{\prime}$) we are at $\tilde{n}^{\prime}=0$ again.
$1\displaystyle 0\leq a,b,c,\tilde{a},\tilde{b},\tilde{c}\leq N-$ 1
$n=a-\tfrac{N}{2}\,,\quad n^{\prime}=b-\tfrac{N}{2}\,,\quad m=c-\tfrac{N}{2}\,,$ (55)
$\tilde{n}=\tilde{a}-\tfrac{N}{2}\,,\quad\tilde{n}^{\prime}=\tilde{b}-\tfrac{N}{2}\,,\quad\tilde{m}=\tilde{c}-\tfrac{N}{2}\,,$ (56)
@@
$\tilde{a}$ $=\phantom{-}a,$ no shift
$\tilde{b}$ $=\phantom{-}a+b-\tfrac{N}{2}$ $=\phantom{-}n+m+\tfrac{N}{2},~{}$ shift by $\textrm{shift by }\tfrac{N}{2}$
$\tilde{c}$ $=-a+b+\tfrac{N}{2}$ $=-n+n^{\prime}+\tfrac{N}{2},~{}$ shift by $\textrm{shift by }\tfrac{N}{2}.$ (57)
Hence, we get the shift vector $\bm{b}=(0,\frac{N}{2},\frac{N}{2})^{\mathrm{T}}$. For example, at $n=n^{\prime}$ we need $\omega=0$ and, thus, $\tilde{c}=\frac{N}{2}$, which is ensured by the shift. This first transformation can be represented by an MPO with $D_{\mathrm{max}}=9$.
For the $\overline{\textit{ph}}$rightarrow pp transformation,
$F_{\textit{pp}}(\nu,\nu^{\prime},\omega)$ $=F_{\overline{\textit{ph}}}(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu)=F_{\overline{\textit{ph}}}(\tilde{\nu},\tilde{\nu}^{\prime},\tilde{\omega}),$ (58)
@@
with $(\tilde{\nu},\tilde{\nu}^{\prime},\tilde{\omega})=(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu)$. Using the same procedure as above, we get the shift vector $\bm{b}=(0,\frac{N}{2}-1,\frac{N}{2})^{\mathrm{T}}$. This affine transformation can be represented by an MPO with $D_{\mathrm{max}}=15$.
@@
The picture corresponds to the following, e.g. for the full vertex
$F_{\textit{ph}}(\nu,\nu^{\prime},\omega)$ $=F_{\textit{pp}}(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}),$ (61a)
$F_{\overline{\textit{pp}}}(\nu,\nu^{\prime},\omega)$ $=F_{\textit{pp}}(\nu,\nu+\omega,\nu^{\prime}-\nu),$ (61b)
$F_{\textit{ph}}(\nu,\nu^{\prime},\omega)$ $=F_{\overline{\textit{pp}}}(\nu,-\nu^{\prime}-\omega,\nu^{\prime}-\nu).$ (61c)
@@ Appendix D Parquet equation in QTT format
Figure 15: (a) Absolute normalized error $||\Delta\Gamma_{d}||:=|\Gamma_{d,\mathrm{parquet}}-\Gamma_{d,\mathrm{exact}}|/||\Gamma_{d,\mathrm{exact}}||_{\infty}$ of the parquet equation with QTTs for $\Gamma_{d}$ compared to the exact values at $\omega=0$ in the fermionic Matsubara frequency plane for $D_{\mathrm{max}}=200$, $R=7$, $\beta=U=1$, $\epsilon=10^{-8}$. Runtime of the parquet equation for $\Gamma_{d}$ for various maximum bond dimensions and grid size parameters $R$ with $\beta=U=1$. The dashed lines indicate (b) the cubic runtime increase with $D_{\mathrm{max}}$ and (c) linear increase by increasing R, which corresponds to an exponential increase in the number of grid points.
@@ Appendix E SDE in QTT format
Figure 16: Plot of the maximum absolute normalized error of the SDE $||\Delta\Sigma||_{\infty}:=||\Sigma_{d,\mathrm{SDE}}-\Sigma_{d,\mathrm{exact}}||_{\infty}/||\Sigma_{d,\mathrm{exact}}||_{\infty}$ with QTTs for the self-energy $\Sigma$ compared to the exact values for various grid sizes depending on the maximum bond dimension with $U=\beta=1$. The dashed lines indicate the results from the dense grid calculation without QTTs.
@@ Appendix F Technical limitations
In the calculations, a specified maximum bond dimension is set not only in the initial TCI, but also in every QTT operation, in order to avoid a blowing up of the bond dimensions since this is the computational bottleneck. As was shown in Sec. V.1, for the initial TCI already bond dimensions slightly above 100 are sufficient to reach maximum normalized errors of $10^{-6}$ of the QTTs with respect to the exact data. However, a problem emerges in the QTT operations, which are at the moment SVD based and make use of the truncate function in ITensors.jl to compress the resulting QTTs back to a certain maximum bond dimension. Furthermore, the fit algorithm used for MPO-MPO contractions relies on the SVD truncation internally [41]. The SVD truncation minimizes the difference between an original MPS and an approximated one in terms of the Frobenius norm. Because the Frobenius norm of the vertex functions grows exponentially with $R$ due to a constant term, the SVD truncation is expected to fail at large $R$; the Frobenius norm reaches $c(2^{R})^{3}\approx c\times 3\times 10^{13}$ at $R=15$ ($c$ is the constant term).
@@
Here, we have observed the truncation error to become more significant the larger the value of the grid parameter $R$ is and even leads to wrong results around $R=15$. In Fig. 17, we show this behavior in case of the full vertex in the density channel $F_{d}$. In Figs. 17(a-b) the absolute normalized error is shown after applying TCI to evaluate a QTT for $F_{d}$ for different maximum bond dimensions. It can be seen that the error in this center ($16\times 16$) fermionic Matsubara frequency box is of $\mathcal{O}(10^{-15})$. Figures 17(c-f) show the error after applying the SVD based truncation to the QTT with maximum bond dimension 160 down to a maximum bond dimension of 140 for various grid sizes determined by $R$. Although using TCI with a maximum bond dimension of 140 [Fig. 17(b)] the QTT was able to reconstruct the exact data with an absolute normalized error of $\mathcal{O}(10^{-15})$, applying the SVD based truncation significantly worsens the results, leading to normalized errors between $10^{-9}$ and $10^{-5}$. Moreover, it can be seen that the error increases significantly with increasing $R$. This is why in the case of the iterative parquet solutions for the Hubbard atom, only results up to $R=11$ are shown, since the resulting maximum normalized error does not improve anymore for larger grids due to the truncation errors. This can also already seen in Fig. 11 for $R=11$, where the maximum normalized error at a maximum bond dimension of 200 is only slightly lower than in the case of $R=10$. However, this should only be a problem of the current implementation and first numerical tests indicate that it is possible to overcome this limitation in the future e.g. by using CI based truncation [34]. This is because the CI-based truncation relies on the maximum norm and thus does not suffer from the divergence of the Frobenius norm.
@@
Figure 17: Absolute normalized error of the reconstructed $F_{d}$ compared to the exact data for the innermost $16\times 16$ fermionic Matsubara frequency grid at $\omega=0,\beta=U=1$. (a) and (b) show the error after using TCI with a set tolerance of $10^{-10}$ and the maximum bond dimensions set to 160 and 140 respectively. (c)–(f) show the error after applying the SVD based truncation to the QTT with maximum bond dimension 160 (a) with a set maximum bond dimension of 140. The truncation error increases with larger values of the grid parameter $R$.